Circular No. 6924

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J2012+381 R. M. Wagner, Ohio State University; S. Starrfield, Arizona State Uni-

versity; and C. Shrader, Goddard Space Flight Center, report: “LONEOS Schmidt camera CCD images obtained on May 27.4 (IAUC 6920) and 28.3 UT by B. Skiff (Lowell Observatory) were examined within the improved RXTE/PCA error circle (IAUC 6922) and compared to the Digital Sky Survey (DSS). No new objects were found in the LONEOS images within the x-ray error circle that are not on the DSS down to mag ∼ 19. Also, none of the objects within the error circle that are present on the LONEOS images appears to have exhibited significant variability with respect to the DSS. In addition, comparison of the LONEOS images within the x-ray error circle does not reveal any new objects on May 28.3 as compared to May 27.4, to a limiting magnitude of ∼ 19. In addition, no objects within the error circle and present on these images exhibit variability exceeding 0.1 mag (3σ), down to mag 18 between the two epochs. Continued monitoring of the error circle is encouraged.”

R. M. Hjellming and M.P. Rupen, National Radio Astronomy Observatory (NRAO); and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “On May 31.25 UT, a radio source was observed with the Very Large Array in the field of the x-ray transient XTE J2012+381 at α = 20^{h}12^{m}37^{s}.70, δ = +38^{o}11′01′′.2 (equinox 2000.0; uncertainty 0′′.5), with flux densities of 2 and 1.5 mJy at 1.4 and 4.9 GHz, respectively. This radio source is well within the error region for the RXTE ASM position reported on IAUC 6920 and 19′′ outside the error region for the RXTE PCA position reported on IAUC 6922. It was not present in an NVSS survey image obtained on 1995 Apr. 28, with an upper limit of ∼ 1 mJy. The source is a candidate for the radio counterpart of XTE J2012+381, but significant radio variability still needs to be established.”

SUPERNOVA 1998S IN NGC 3877 Additional CCD photometry by D. Hanžl, Brno, Czech Republic (cf.

IAUC 6885): Apr. 27.971 UT, V = 13.57 ± 0.02, B − V = +1.01 ± 0.04; 30.838, 13.66 ± 0.02, +0.95 ± 0.03; May 2.925, 13.77 ± 0.04, +0.98 ± 0.09; 5.904, 13.88 ± 0.02, +1.00 ± 0.06; 8.887, 13.97 ± 0.01, +0.97 ± 0.03; 10.919, 14.07 ± 0.02, +1.03 ± 0.03; 20.876, 14.76 ± 0.03, –; 28.896, 15.08 ± 0.05, –; 29.878, 15.1 ± 0.1, –; 31.926, 15.12 ± 0.04, –.

1998 June 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green